OpenStudy (anonymous):

Which of the following is an irrational number? A. B. 1/2 C. D. 3.8

OpenStudy (mertsj):

None of those

OpenStudy (anonymous):

is it 18 or 49 ?

OpenStudy (mertsj):

An irrational number is like this: \[\sqrt{2}, \pi, e \sqrt{15}, etc\]

OpenStudy (anonymous):

ohh ty

OpenStudy (mertsj):

yw